OpenStudy (anonymous):

Can someone explain to me how to do this!?! I dont understand at all :(. Write the standard form of the equation of the line passing through the point (2, 4) and perpendicular to the line - 4x - y = 5. [A] - x + 4y = –14 [B] x - 4y = –14 [C] - 4x + y = –4 [D] none of these

OpenStudy (anonymous):

given equation -4x-y=5 which is perpendicular to the points (2,4), find the equation of the line? ------------------------------------------------- (1) find the slope -4x-y=5 -y=4x+5 y=-4x+5 thus, m=-4 since it is perpendicular to the ask equation of the line, find the inverse of that slope(m). (2) //m=1/4 (3) using the //m and the given point (2,4), find the equation of the line. . using point-slope form: \[(y-y _{1})=m(x-x _{1})\]substitute. . . (y-4)=1/4(x-2) y-4 =1/4x-1/2 (y-4= 1/4x-1/2)4 4y-16=x-2 -x+4y=14 or x-4y=-14